Obtain an equation of a line with point -3,-2 & 1,7.show solution pls.thank you
Answer by nerdybill(7384)

Obtain an equation of a line with point -3,-2 & 1,7
.
First, determine slope between the two points:
m(slope) = (y2 - y1)/(x2 - x1)
m(slope) = (7 - (-3))/(1 - (-3))
m(slope) = (7 + 3)/(1 + 3)
m(slope) = 10/4
m(slope) = 5/2
.
Using the slope (5/2) and one point (1,7) plug into "point-slope" form:
y - y1 = m(x - x1)
y - 7 = (5/2)(x - 1)
y - 7 = (5/2)x - 5/2
y = (5/2)x - 5/2 + 7
y = (5/2)x - 5/2 + 14/2
y = (5/2)x + 9/2 (this is what they're looking for in "slope-intercept" form)